# Learning Interior Point Method for AC and DC Optimal Power Flow

### v1

@@ Abstract
This paper proposes a feasibility-guaranteed learning interior point method (L-IPM) to solve both AC and DC optimal power flow (OPF) problems. Given the criticality of OPF, the proposed L-IPM uses a hybrid learning-model approach rather than relying solely on a simple black-box prediction. The traditional IPM follows a central path from an initial point to the optimal solution. However, each iteration involves solving large linear systems, which becomes increasingly expensive as the matrices grow more ill-conditioned in later steps. To address this, we model the IPM trajectory as a time series and train a Long Short-Term Memory (LSTM) network to project the IPM central path using only the first few stable iterations, which carry the most informative features about the path to optimality. We introduce a grid-informed methodology that enforces operational constraints on generation, voltage magnitudes, and line flows to ensure feasibility. The grid-informed LSTM serves as a tool for the IPM central path projection and, followed by a final IPM refinement step, significantly reduces the total number of iterations and time required for convergence. We use a sampling method to generate a wide range of load scenarios to improve generalization across diverse operating conditions, efficiently covering the power system’s operational space. Simulation results on a 2869-bus European high-voltage transmission system show that the proposed L-IPM significantly reduces solution time by up to 94%, while maintaining accuracy and feasibility of the solution. By leveraging early iterations and bypassing the final ill-conditioned and computationally demanding steps of traditional IPM, the proposed L-IPM reduces the number of required iterations by up to 85.5%. Since solution feasibility is also guaranteed, L-IPM outperforms the conventional IPM in both computational efficiency and robustness.
Index Terms:
Learning interior point method, grid-informed LSTM, AC and DC optimal power flow
@@ I Introduction
Several methods have been proposed to accelerate OPF computation and improve its efficiency, ranging from advanced mathematical formulations to emerging paradigms such as quantum computing [4, 5, 6]. Among recent algorithmic advancements, techniques such as semidefinite programming [7], quantum-enhanced optimization [8], sequential quadratic programming [9], and methods based on the augmented Lagrangian frameworks [10] have shown considerable promise. In addition to these optimization-based approaches, some researchers have explored hardware and parallelization enhancements, such as GPU acceleration [11] and high-performance computing [12], which can improve performance but require access to specialized infrastructure. Others have used heuristic and metaheuristic algorithms [13], which may offer faster solutions but do not guarantee global optimality.
@@
More recently, attention has turned toward machine learning-based methods, including grid-informed neural networks, supervised learning approximations, reinforcement learning, and graph neural networks [14]. These approaches are typically reduce problem dimensionality or approximate OPF solutions in a black-box manner, often without modifying the internal structure of the OPF or optimization algorithm. While promising, these methods do not always guarantee consistent feasibility across varying system conditions.
@@
Among all OPF solution methods, the interior point method (IPM) remains one of the most popular and reliable techniques [15]. IPM operates by iteratively applying Newton’s steps to optimize system variables, where each step involves solving a large linear system. Typically, the most computationally expensive part of the algorithm is the solution of these linear systems [16]. Some researchers have explored quantum linear solvers to address this computational burden, which offer potential asymptotic speedups [17]. However, hardware limits, errors, and scalability challenges still prevent their use in large-scale OPF [18]. Other approaches to enhance IPM include heuristic methods such as inexact Newton strategies [19] and machine learning-based techniques [14]. Machine learning has primarily been used to provide warm starts for IPM by predicting good initial guesses. While helpful in some cases, these methods do not modify the IPM itself, and as a result, even high-quality warm starts do not consistently lead to reduced IPM iterations and computation time. We want to use machine learning directly to enhance the IPM algorithm’s internal behavior or to learn from its iterative patterns. If the number of Newton steps required for convergence could be reduced through such learning, the overall computational complexity of IPM could be enhanced.
@@
IPM follows a trajectory toward the optimal solution, which can be interpreted as a time series, as it progresses iteratively toward convergence. In this work, we propose to model the trajectory of the central path as a data sequence and learn its behavior using a Long Short-Term Memory (LSTM) network. A key observation motivating this approach is that the most informative patterns in the IPM trajectory typically emerge during the early iterations, while the final iterations primarily serve to refine feasibility and enforce constraint satisfaction. By leveraging this property, we can apply path sampling techniques to project the optimal solution, enabling a significant reduction in the number of Newton steps required and, consequently, a substantial decrease in the overall OPF computation time.
@@
Applying LSTM to project the central path of IPM presents several challenges. Since OPF is a constrained optimization problem, direct projections do not inherently guarantee feasibility. Infeasible solutions can lead to constraint violations, which compromise system reliability and may have catastrophic consequences in real-world power grid operations. To address this, we propose a grid-informed LSTM (GI-LSTM) that incorporates system constraints into the learning process. Another key challenge lies in generating a sufficiently diverse and representative training dataset. For a given power network, the number of possible operating conditions grows exponentially with the number of buses and load variations. Ensuring the generalization and reliability of the learning IPM (L-IPM) approach, therefore, requires careful attention to both data generation and model architecture.
@@
  • A novel L-IPM framework is proposed for solving the OPF problem, designed to address these computational and modeling challenges. Our analysis reveals that the early IPM iteration typically corresponds to a system with the lowest condition number. In contrast, subsequent iterations become increasingly ill-conditioned, leading to low accuracy of the solution. Motivated by this insight, we use only the initial few iterations as inputs to the GI-LSTM model.
  • To ensure feasibility, we incorporate a grid-informed loss function that penalizes violations of generator limits, voltage magnitudes, and power flow constraints.
  • To efficiently capture a wide range of system conditions, we employ Latin hypercube sampling (LHS) to generate diverse load scenarios without the need for exhaustive enumeration [20].
  • The GI-LSTM projection is refined using the classic IPM to ensure convergence to an exact and feasible solution.
Simulation results on various systems, including a realistic large 2869-bus system, show that the proposed L-IPM significantly reduces the number of IPM iterations, Newton steps, and solution time while maintaining both solution accuracy and OPF feasibility.
@@ II-A OPF Formulation
where $p_{i}$ is the real power generation at bus $i$, for all $i\in\mathbb{G}$, and $a_{i},b_{i},c_{i}$ are cost coefficients. OPF is subject to AC power flow equations that enforce active and reactive power balance at each bus:
@@
where $p_{i}$ and $q_{i}$ are the real and reactive power injections at bus $i$, $v_{i}$ is the voltage at bus $i$, and $v_{j}^{*}$ is the complex conjugate of the voltage at bus $j$. The problem is further constrained by generator limits in (3), voltage constraints 4, and transmission line thermal limits in (5).
@@ II-B Classic IPM
$-\mu\sum_{i\in\mathbb{G}}\left(\log s_{i}^{+}+\log s_{i}^{-}\right)-\mu\sum_{i\in\mathbb{B}}\left(\log v_{i}^{+}+\log v_{i}^{-}\right)$ (6)
The term $\Re$ ensures the Lagrangian remains real-valued. To solve the Lagrangian formulation, we construct the Karush-Kuhn-Tucker (KKT) conditions, which are derived with respect to all primal variables (e.g., $\mathcal{PD}=\left\{p_{i},\ q_{i},\ v_{i},\ \theta_{i}\right\}$), slack variables (e.g., $\mathcal{SD}=\left\{s_{i}^{\pm},\ v_{i}^{\pm},\ \sigma_{ij}\right\}$), and dual variables associated with the constraints. Newton’s method is then applied to solve this system iteratively. At each iteration, a linear system is formed with a coefficient matrix known as the KKT matrix.
@@ III Proposed Grid-Informed L-IPM
The IPM-based OPF follows a structured trajectory known as the central path, illustrated in Fig. 1. We have observed that the initial iterations of the central path contain valuable information about its trajectory and, consequently, the final optimization solution. These early steps typically cover a significant portion of the path toward optimality. One can also clearly recognize that the IPM central path closely resembles a time series. Given this characteristic, LSTM networks are well-suited for leveraging early iteration data to predict the solution.
@@
Figure 1: IPM path to the optimal solution.
@@ III-A Learning Model Architecture
TABLE I: Dataset Structure for GI-LSTM Training
T Step Bus ID Load (MW) Gen (MW) Volt. (p.u.) Ang. (°) Feas.
1 1 100.5 120.3 1.02 3.5 Yes
2 1 101.8 119.7 1.01 3.6 Yes
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
@@ III-B LSTM Training Mapping Strategy
Given a set of system conditions, the LSTM model is trained to learn a temporal mapping function from a sequence of past states to the final OPF solution [22]. Formally, the model learns:
@@ III-C Grid-Informed LSTM for OPF Feasibility
The LSTM network projects the final iteration of the IPM, which is expected to lie close to the optimal OPF solution. This is illustrated in Fig. 4, which provides a zoomed-in view of the optimal solution shown in Fig. 2(a). While LSTM is effective in learning the convergence trajectory of IPM, ensuring that the LSTM projection of the IPM central path remains feasible or stays within the green region surrounding the optimal solution is critical for maintaining OPF feasibility and satisfying power system operational constraints. If the projection falls within the red zone, it would lie outside the OPF feasible region, thus violating the operational constraints. This is problematic because, as discussed later, the proposed L-IPM approach relies on a hybrid learning-model framework that uses the learner’s projection to guide IPM refinement near the optimal solution. If the learner’s projection is infeasible and falls outside the feasible region, the IPM may struggle to refine it and converge to the true OPF optimal solution.
@@
where the first two terms ensure accurate predictions for generator values and bus voltage magnitudes, while the third penalizes constraint violations in power flows. The hyperparameter $\omega$ balances accuracy and feasibility. The first two terms $\omega_{1}$ and $\omega_{2}$ ensure accurate predictions of generator outputs and voltage magnitudes, while the third penalizes power flow constraint violations. Among these, $\omega_{3}$ is typically assigned the highest value to strongly discourage infeasible solutions.
@@ III-D2 Regularization Techniques
To prevent overfitting and enhance model generalization across diverse OPF scenarios, regularization strategies are used. Dropout regularization randomly deactivates neurons during training, with an appropriate dropout rate selected to balance robustness and information retention. Additionally, L2 regularization (weight decay) penalizes large weights, simplifying the model to improve generalization [22].
@@ III-E L-IPM for OPF Solution
Reducing the number of IPM iterations before predicting the final solution allows for substantial computational savings. However, it is important to ensure that the GI-LSTM model generalizes well to diverse OPF conditions. Using L-IPM, which combines GI-LSTM predictions with IPM refinement, we ensure the feasibility and accuracy of the final solution. The process begins by training GI-LSTM on historical OPF solutions obtained from IPM. Once trained, GI-LSTM generates an initial guess $y_{t}^{LSTM}$ for the OPF solution, which is then refined using a final IPM optimization step:
@@
The proposed L-IPM framework modifies the classic IPM algorithm to achieve faster convergence toward the same solution. Feasibility and accuracy are preserved through a final IPM refinement step, as illustrated in Fig. 3.
@@ V-A Simulation Setting
To evaluate the effectiveness and scalability of the proposed L-IPM, we conducted tests on power systems of varying sizes. Our analysis began with a small-scale 3-bus system and a standard 24-bus system, followed by larger systems comprising 118 and 2869 buses to assess the scalability of the method. The GI-LSTM model was implemented in Python, while the IPM-based OPF simulations were carried out using the MATPOWER 8.0 toolbox in MATLAB.
@@ V-B Look-Back Window Determination
To determine the appropriate number of early iterations to input into the GI-LSTM, we conducted multiple simulations. A single iteration does not provide enough historical context for the GI-LSTM to capture temporal patterns and interdependencies. Therefore, we increased the number of early iterations, or the look-back window, from one to five and observed that in many cases, the output became more accurate with a higher number of iterations. However, there were instances where increasing the number of iterations led to a slight decrease in accuracy. In Fig. 5, we compare the $R^{2}$ scores for two systems. We observe that setting the look-back to three provides satisfactory performance, and for the 2869-bus system, it achieves the best performance overall. Moreover, in many cases, increasing the look-back from three to four or five does not significantly improve precision. Based on these observations, we selected a look-back of three iterations to feed into the GI-LSTM.

### v2

# Learning Interior Point Method Central Path Projection for Optimal Power Flow

@@ Abstract
This paper proposes a learning-based approach to accelerate the interior-point method (IPM) for solving optimal power flow (OPF) problems by learning the structure of the IPM central path from its early stable iterations. Unlike traditional learning models that attempt to predict the OPF solution directly, our approach learns the structure of the IPM trajectory itself, since even accurate predictions may not reliably reduce IPM iterations. The IPM follows a central path that iteratively progresses toward the optimal solution. While this trajectory encodes critical information about the optimization landscape, the later iterations become increasingly expensive due to ill-conditioned linear systems. Our analysis of the IPM central path reveals that its initial segments contain the most informative features for guiding the trajectory toward optimality. Leveraging this insight, we model the central path as a time series and use a Long Short-Term Memory (LSTM) network to project the path using only the first few stable iterations. To ensure that the learned trajectory remains within the feasible region—especially near the optimal point—we introduce a grid-informed mechanism into the LSTM that enforces key operational constraints on generation, voltage magnitudes, and line flows. This framework, referred to as Learning-IPM (L-IPM), significantly reduces both the number of IPM iterations and overall solution time. To improve generalization, we use a sampling-based strategy to generate a diverse set of load conditions that effectively span the operational space. Simulation results across a range of test systems—including a 2869-bus European transmission network—demonstrate that L-IPM achieves up to a 94% reduction in solution time and an 85.5% reduction in iterations, without compromising feasibility or accuracy. We also compare L-IPM with classical solution-prediction-based warm starts and show that our central-path projection strategy offers superior performance.
@@ I-A Background
Several methods have been proposed to accelerate OPF computation and improve its efficiency, ranging from advanced mathematical formulations to emerging paradigms such as quantum computing [17, 7]. Among recent algorithmic advancements, techniques such as semidefinite programming [24], quantum-enhanced optimization [5], sequential quadratic programming [22], and methods based on the augmented Lagrangian frameworks [13] have shown considerable promise. In addition to these optimization-based approaches, some researchers have explored hardware and parallelization enhancements, such as GPU acceleration and high-performance computing [2], which can improve performance but require access to specialized infrastructure. Others have used heuristic and metaheuristic algorithms [8], which may offer faster solutions but do not guarantee global optimality.
@@
More recently, attention has turned toward machine learning-based methods, including grid-informed neural networks, supervised learning approximations, reinforcement learning, and graph neural networks [15, 4]. Many of these approaches typically reduce problem dimensionality or directly predict the OPF solutions. As a result, they generally provide less direct leverage over the optimizer’s iterative behavior.
I-B Motivation and Research Gap
@@
Among all OPF solution methods, the interior point method (IPM) remains one of the most popular and reliable techniques [12]. IPM operates by iteratively applying Newton’s steps to optimize system variables, where each step involves solving a large linear system of equations. Typically, the most computationally expensive part of the algorithm is solving these linear systems [10]. Some researchers have explored quantum linear solvers to address this computational burden, which offer potential asymptotic speedups [6]. However, hardware limits, errors, and scalability challenges still prevent their use in large-scale OPF [14]. Other approaches to enhance IPM include heuristic methods such as inexact Newton strategies [23] and machine learning-based techniques [15].
Existing machine learning-assisted OPF efforts can be broadly grouped into three categories: (i) direct end-to-end prediction of the solution variables, which may produce erroneous outputs with limited control over feasibility and solution quality; (ii) warm-start approaches that predict high-quality solution guesses to start closer to the optimum, yet for iterative solvers such as IPM, this does not reliably reduce iterations or runtime; and (iii) iteration-reduction or feasibility-oriented schemes that steer iterates toward feasible regions, often improving feasibility but potentially sacrificing optimality [20, 9, 21]. Among these, warm-start strategies are especially appealing in practice because they can provide accurate initial primal points; however, even accurate initial primal points do not reliably translate into fewer Newton steps or lower runtime. This is because IPM convergence depends not only on the primal variables, but also on the consistent initialization and subsequent evolution of dual variables, slack variables, and barrier/centrality conditions that maintain feasibility and guide the iterates along the central path [23]. In practice, even when IPM starts from a nearly optimal primal solution, it may still need about the same number of iterations, or even more, to restore centrality and satisfy the full Karush–Kuhn–Tucker (KKT) conditions. Case studies show that an optimal OPF warm start does not always reduce the iteration count; in some cases, IPM converges in fewer iterations without it. In contrast, our objective is to reduce runtime by directly reducing the number of IPM iterations while preserving feasibility. Rather than focusing solely on the initial point, we exploit the iterative behavior of the IPM central path to learn informative update directions and accelerate convergence.
@@
IPM central path follows a trajectory toward the optimal solution, which can be interpreted as a time series, as it progresses iteratively toward convergence. In this work, we propose modeling the trajectory of the central path as a data sequence and learning its behavior using a Long Short-Term Memory (LSTM) network, owing to its ability to capture temporal dependencies in sequential data [16]. A key characteristic motivating this approach is that the most informative patterns in the IPM central path trajectory emerge during the early iterations, while the final iterations primarily serve to guarantee feasibility and enforce constraint satisfaction. By leveraging this property, we can apply path sampling techniques to project the IPM central path, enabling a significant reduction in the number of Newton steps required and, consequently, a substantial decrease in the overall OPF computation time.
I-C Contributions
@@
Applying LSTM to project the central path of IPM presents several challenges. Since OPF is a constrained optimization problem, direct projections without attention to the grid physics and optimization constraints do not inherently guarantee a high-quality solution. Infeasible solutions can lead to constraint violations, compromising system reliability and potentially having catastrophic consequences in real-world power grid operations. To address this, we propose a grid-informed LSTM (GI-LSTM) that incorporates system constraints into the learning process. Another key challenge lies in generating a sufficiently diverse and representative training dataset. For a given power network, the number of possible operating conditions grows exponentially with the number of buses and load variations. Ensuring the generalization and reliability of the learning IPM (L-IPM) approach, therefore, requires careful attention to both data generation and model architecture.
@@
  • The IPM central-path trajectory for OPF is analyzed to extract informative patterns from the early iterations, which capture valuable relationships between grid physics and optimization variables. Building on these insights, we propose an L-IPM framework that leverages early-iteration behavior to project the central path toward a high-quality OPF solution.
  • To address OPF computational and modeling challenges, the proposed approach reduces the required number of IPM iterations by relying on the early steps, where the underlying matrices are best conditioned and most reliable to compute. In contrast, later iterations tend to become increasingly ill-conditioned, which can degrade numerical reliability and increase computational burden.
  • To ensure the solution quality, we incorporate a grid-informed loss function that penalizes violations of generator limits, voltage magnitudes, and power flow constraints.
  • To capture a wide range of system conditions, we use Latin hypercube sampling (LHS) to generate diverse load scenarios without the need for exhaustive enumeration [11].
Simulations on various systems, including a realistic large 2869-bus system, show that the proposed L-IPM significantly reduces the number of IPM iterations, Newton steps, and solution time while maintaining OPF solution accuracy.
@@ II-A OPF Formulation
OPF is subject to AC power flow equations that enforce active and reactive power balance at each bus:
@@
The problem is further constrained by generator, voltage, and transmission line thermal limits in (3), 4, and (5).
@@ II-B Classic IPM
$-\mu\sum_{i\in\mathbb{G}}\left(\log s_{i}^{+}+\log s_{i}^{-}\right)-\mu\sum_{i\in\mathbb{B}}\left(\log v_{i}^{+}+\log v_{i}^{-}\right)$
The term $\Re$ ensures the Lagrangian remains real-valued. To solve the Lagrangian formulation, we construct the KKT conditions, which are derived with respect to all primal variables (e.g., $\mathcal{PD}=\left\{p_{i},\ q_{i},\ v_{i},\ \theta_{i}\right\}$), slack variables (e.g., $\mathcal{SD}=\left\{s_{i}^{\pm},\ v_{i}^{\pm},\ \sigma_{ij}\right\}$), and dual variables associated with the constraints. Newton’s method is then applied to solve this system iteratively. At each iteration, a linear system is formed with a coefficient matrix known as the KKT matrix.
@@ III Proposed Grid-Informed L-IPM
The IPM-based OPF follows a structured trajectory known as the central path, illustrated in Fig. 1. The initial iterations of the central path contain information about its trajectory and, consequently, the final optimization solution. These early steps typically cover a significant portion of the path toward optimality. The IPM central path closely resembles a time series. Given this characteristic, LSTM networks are well-suited for leveraging early iteration data to predict the solution.
@@
Figure 1: IPM and the proposed L-IPM paths to the optimal solution.
@@ III-A Learning Model Architecture
TABLE I: GI-LSTM Training Data: Primal–Dual IPM Trajectory (Per Scenario $s$ and iteration $k$)
$s$ Load Primal $\mathbf{x}_{k}$ Eq. dual $\boldsymbol{\lambda}_{k}$ Ineq. dual $\boldsymbol{\mu}_{k}$ Feas.
11 $\mathbf{d}^{(1)}$ $P_{G},V,\theta,\ldots]$ $\lambda^{\text{Pbal}},\lambda^{\text{Qbal}},\ldots$ $\mu^{\text{line}},\mu^{V},\mu^{P_{G}},\ldots$ $(\text{Yes})$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
@@ III-B LSTM Training Mapping Strategy
Given a set of system conditions, LSTM is trained to learn a temporal mapping function from a sequence of past states to the final OPF solution [16]. Formally, the model learns:
@@ III-C Grid-Informed LSTM for OPF Feasibility
The LSTM network projects the final iteration of the IPM or the optimal OPF solution. This is illustrated in Fig. 4, which provides a zoomed-in view of the optimal solution shown in Fig. 2(a). While LSTM is effective in learning the convergence trajectory of IPM, ensuring that the LSTM projection of the IPM central path remains feasible or stays within the green region surrounding the optimal solution is critical for maintaining OPF feasibility and satisfying power system operational constraints. If the projection falls within the red zone, it would lie outside the OPF feasible region, thus violating the operational constraints. This is problematic because the proposed L-IPM approach relies on a learning-model framework that uses the learner’s central path projection to find the optimal solution. If the learner’s projection is infeasible and falls outside the feasible region, the last IPM step does not verify the optimality of the solution.
@@
Hyperparameters $\omega_{1}$ to $\omega_{3}$ balance accuracy and feasibility. The first two terms $\omega_{1}$ and $\omega_{2}$ ensure accurate predictions of generator outputs and voltage magnitudes, while the third penalizes power flow constraint violations. Among these, $\omega_{3}$ is typically assigned the highest value to strongly discourage infeasible solutions.
@@ III-D2 Regularization Techniques
To prevent overfitting and enhance generalization across diverse OPF scenarios, we use dropout and L2 regularization. Dropout randomly deactivates neurons during training, with the rate tuned to balance robustness and information retention, while L2 (weight decay) penalizes large weights to simplify the model and improve generalization [16, 3].
@@ III-E L-IPM for OPF Solution
Reducing the number of IPM iterations before predicting the optimal solution allows for substantial computational savings. However, it is important to ensure that the GI-LSTM model generalizes well to diverse OPF conditions. Using L-IPM, which combines GI-LSTM predictions with IPM refinement, we enforce the feasibility and accuracy of the central path projection and, consequently, the OPF solution. The process begins by training GI-LSTM on historical OPF solutions obtained from IPM. Once trained, GI-LSTM generates the central path projection $y_{t}^{LSTM}$ for the OPF solution, which is then the optimality is verified using a final IPM optimization step:
@@
The proposed L-IPM framework modifies the classic IPM algorithm to achieve faster convergence toward the same solution. Feasibility and accuracy are enforced through GI-LSTM and the final IPM verification steps, as illustrated in Fig. 3.
@@ V Numerical Results
We evaluate the approach on a diverse set of test systems, ranging from small networks to large-scale, real-world cases, and examine its behavior under varying operating conditions. To provide insight into robustness, we consider a wide range of load scenarios and analyze feasibility/infeasibility detection and performance trends across these conditions. For comparison, we also report the convergence behavior of a classical warm-start prediction strategy, highlighting key differences in iteration dynamics and convergence behavior.
@@ V-A Simulation Setting
To evaluate the effectiveness and scalability of the proposed L-IPM, we conducted tests on power systems of varying sizes. Our analysis began with a small-scale 3-bus system and a standard 24-bus system, followed by larger systems comprising 118 and 2869 buses to assess the scalability of the method. The GI-LSTM model was implemented in Python using TensorFlow, which also enabled the implementation of the proposed grid-informed loss function by augmenting the standard prediction error with differentiable feasibility-penalty terms. The IPM-based OPF simulations were carried out using the MATPOWER 8.0 toolbox in MATLAB.
@@ V-B Look-Back Window Determination
To determine the appropriate number of early iterations to input into the GI-LSTM, we conducted multiple simulations. A single iteration does not provide enough historical context for the GI-LSTM to capture temporal patterns and interdependencies. Therefore, we increased the number of early iterations, or the look-back window, from one to five and observed that in many cases, the output became more accurate with a higher number of iterations. However, there were instances where increasing the number of iterations led to a slight decrease in accuracy. Fig. 5 reports the $R^{2}$ scores for two representative systems for clarity. We observe that setting the look-back to three provides satisfactory performance, and for the 2869-bus system, it achieves the best performance overall. Moreover, in many cases, increasing the look-back from three to four or five does not significantly improve precision. Based on these observations, we selected a look-back of three iterations to feed into the GI-LSTM.